OpenStudy (anonymous):

log[8](1/8)=_log[8](2)

OpenStudy (mertsj):

log (base 8) of 1/8 = -1 log (base 8) of 2 = 1/3

OpenStudy (anonymous):

so its -1log[8](2)???

OpenStudy (mertsj):

No. substitute. -1= x(1/3)